### v1

# Scene-Specific Trajectory Sets: Maximizing Representation in Motion Forecasting

Abstract
Representing diverse and plausible future trajectories of actors is crucial for motion forecasting in autonomous driving. However, efficiently capturing the true trajectory distribution with a compact set is challenging. In this work, we propose a novel approach for generating scene-specific trajectory sets that better represent the diversity and admissibility of future actor behavior. Our method constructs multiple trajectory sets tailored to different scene contexts, such as intersections and non-intersections, by leveraging map information and actor dynamics. We introduce a deterministic goal sampling algorithm that identifies relevant map regions and generates trajectories conditioned on the scene layout. Furthermore, we empirically investigate various sampling strategies and set sizes to optimize the trade-off between coverage and diversity. Experiments on the Argoverse 2 dataset demonstrate that our scene-specific sets achieve higher plausibility while maintaining diversity compared to traditional single-set approaches. The proposed Recursive In-Distribution Subsampling (RIDS) method effectively condenses the representation space and outperforms metric-driven sampling in terms of trajectory admissibility. Our work highlights the benefits of scene-aware trajectory set generation for capturing the complex and heterogeneous nature of actor behavior in real-world driving scenarios.
Index Terms:
motion forecasting, planning, safety, trajectory sets, plausibility
I Introduction
In Autonomous Driving, motion forecasting of actors surrounding the ego vehicle has been a persistent challenge due to the evolving complexities associated with a scene, exacerbated by the inherent nature of dynamic interactions between different traffic participants. Capturing such interactions through discrete intent has been successful to some extent, but such representations lack multi-modality to capture the true posterior distribution of the actors. Consequently, motion forecasting demands multi-modality of plausible outcomes, defining the intended actions required to reach a particular goal position [1]. The associated model uncertainty can result in system behavior that is non-conformant with both physics and scene geometry. In certain situations, this non-conformance can lead to fatal consequences. [2]
Trajectory sets play a crucial role in encoding physics into the motion forecasting layer of an autonomous driving (AD) stack, as demonstrated in the works of [3, 4, 5]. These sets provide a structured representation of the possible future trajectories of traffic participants sampled from the training distribution. By incorporating such trajectory sets into the motion forecasting process, the AD system can generate predictions that are more consistent with the vehicle kinematics and geometric constraints of the scene.
Figure 1: Trajectory set, which represents the possible full actor representation within the scene
Existing methods compute sets independent of the underlying scene representation (acquired from HD maps) and solely rely on minimizing a cost function based on some metrics. This leads to pertinent gaps about constructing an effective set, whereby the cost function includes a geometric correlation between reachability and the implied metrics. Sets, therefore, must be carefully designed to balance the diversity of plausible outcomes; failure to achieve this balance can lead to suboptimal system behavior arising out of sparsity problems within the output representation space [5]. This is because constructing such an optimal set is a max-min k-dispersion problem, which belongs to the NP-hardness complexity [6]. Through this work, our contributions are twofolds;
  • Scene-Specific Trajectory Set Generation: We introduce a novel algorithm for creating scene-specific trajectory sets by leveraging HD map information and actor state. Our method constructs a directed graph from the vector map, performs a depth-first search to identify goal lanes, and categorizes scenes into intersection and non-intersection bins. This approach enables the generation of tailored trajectory sets that better represent the geometric constraints and likely actor behaviors for each scene type, improving both diversity and plausibility of trajectories.
  • Empirical Analysis of Sampling Methods and Set Sizes: Through extensive experiments on the Argoverse 2 dataset, we investigate the impact of different sampling strategies and set sizes on the diversity and plausibility of the generated trajectory sets. We conduct a comprehensive empirical study comparing metric-driven, random sampling, and our proposed Recursive In-Distribution Subsampling (RIDS) methods across various set sizes (1k to 5k trajectories). The results demonstrate that sampling from a uniform distribution using RIDS outperforms the metric-driven method in terms of retaining more plausible trajectories while maintaining comparable displacement errors. We empirically show that RIDS can effectively condense the full representation space to a manageable set size while preserving coverage and plausibility. Furthermore, our analysis reveals the optimal balance between set size and performance metrics, providing insights for efficient trajectory set generation in downstream motion forecasting tasks.
II Related Works
The trajectory prediction problem has been predominantly addressed using regression approaches, which aim to predict a set of waypoints that closely resemble the ground truth trajectory. During training, the loss between the ground truth and the model prediction is optimized. However, regression methods often rely on anchors to enforce diversity between trajectories and require anchors for preventing the models against mode collapse [7, 8]. Regression-based approaches without explicit integration of prior knowledge leads to predictions which go off-road [9] without respecting the neither the underlying kinematics nor the environment [10, 5].
Trajectory Sets in Planning and Prediction Trajectory sets have emerged as a choice in motion prediction, particularly in autonomous driving, as they model the uncertain future as a distribution over a compact set of possible trajectories for each agent. Predominantly, trajectory sets or trim trajectories have been used in planning [11] to resolve for a valid path at a constant time and memory. This flexibility has demonstrated state-of-the-art performance [3, 12, 4] by reformulating the problem as a classification where each of the trajectories in the set represents an individual anchor [8]. This effectively avoids mode collapse, which can occur with models when trained with a regression loss. While traditional methods like Bayesian approaches [13] can be computationally intensive and may pose limitations for real-world applications, whereas trajectory sets offer a practical and empirically superior alternative.
Multiple Sets for Improved Representation Single-set representations often fail to capture the true posterior distribution of ground truth probabilities in real-world scenarios due to sparsity and model expressivity issues [5, 14]. We propose using multiple subsets to create a denser representation space that better approximates ground truth probabilities and trajectory existence. The effectiveness of these sets depends on scene clustering methods, which can be supervised or semi-supervised [15]. Various approaches exist, from using traffic actors’ state vectors with criticality metrics [16, 17, 18] to clustering based on agent trajectories [19]. Our method employs an L-norm metric between lane centerlines and the target actor’s current state, avoiding the look-ahead bias present in other methods that require full actor state representation to create cluster bins.
In our previous work [5], we investigated the approach of pruning trajectory sets based on their intersection with non-drivable regions. However, the generated set lacked priors knowledge with map information, leading to a significant number of trajectories being discarded due to non-plausible states. To address this limitation and build upon the new concept of multiple sets, we propose an alternative dataset-agnostic methodology that understands map information and produces multiple sets, helping to overcome sparsity issues and maintain an optimal balance between diversity and admissibility, as discussed by Park et al. [20].
III Problem Formulation
We begin by establishing the formal notations that will be consistently utilized throughout this document. Consider a full set of trajectories captured from the training distribution as $T_{full}$. Our objective is to deduce a subset $T^{\prime}\subset T_{full}$ that maximizes the diversity and plausibility of trajectories for a cluster of scenes $M_{i}$.
$M=\{M_{i}\mid i\in\mathbb{Z}^{+},1\leq i\leq n\}$ (1)
Let $M$ be the set of all scene clusters where $M_{i}$ represents an individual cluster of scenes and $i$ is a positive integer (denoted by $\mathbb{Z}^{+}$) with $n$ being the total number of clusters. $|T^{\prime}|$ denotes the cardinality of a set $T^{\prime}$. A trajectory $T_{i}$ of an actor $a_{i}$ is represented by a sequence of states, as shown in eq. 2, where $s_{i}^{t}$ represents the center point location $(x,y)$ of $a_{i}$ at a time $t$ in a Cartesian coordinate frame.
$T_{i}=\left\{\underbrace{s_{i}^{-t_{obs}},s_{i}^{-t_{obs}+1},s_{i}^{-t_{obs}+2},...s_{i}^{0}}_{\text{Past states}},\underbrace{s_{i}^{1},s_{i}^{2},...,s_{i}^{t_{fut}}}_{\text{Future states}}\right\}$ (2)
Here, $t_{\text{obs}}$ represents the end of the observation window, and $t_{fut}$ represents the end of the prediction horizon. This notation aligns with the formalism used throughout the paper, particularly focusing on the representation of trajectories as sequences of states over time.
While promoting diversity ($D$) among the trajectories within $T^{\prime}$, it is important to note that coverage alone is insufficient to ensure accurate and plausible output. $T^{\prime}$ should also be realistic or plausible ($P$), and therefore should not contain states that are physically infeasible to reach for a given scene $S$ in $M_{i}$. Additionally, $S$ includes lane information that delineates the probable motion paths an actor can choose to reach a goal position. Within this work, a Focal Track ($FT$) is of an object type car.
$\max_{T^{\prime}\subset T_{full}}\left[D(T^{\prime})+P(T^{\prime}|M_{i})\right]$ (3)
Thus, we aim to maximize eq. 3 where we search for a set which not only optimizes for the coverage metric but also be diverse enough to sustain the environmental constraints without getting pruned [5]; a balance between diversity and plausibility.
Figure 2: Candidate set for intersection scenes
Figure 3: Candidate set for non-intersection scenes
IV Scene-Specific Set Generation
Figure 4: Past trajectory is shown along with the Future trajectory of an FT. The Grey edges represent the connectivity between different lanes with blue nodes representing the source nodes and Violet denotes the possible goal lanes which the FT can move along
IV-A Algorithm
The deterministic goal sampler operates on the scene $S$ and the heading angle $\theta$ w.r.t the Focal Track to generate a set of $K$ goal lanes, denoted as $L_{goal}$. This operation can be succinctly represented by the following eq. 4:
$L_{goal}=f(S,\theta)$ (4)
Given a vector map of $S$ and the state information of $FT$, the algorithm constructs a directed graph $G$ from the map. The last observation state ($s_{i}^{0}$) of the $FT$ is used to find the nearest nodes by filtering for edges with parallel direction vectors to that of the $FT$. This helps us find for the edges which have a higher similarity value w.r.t the $FT$’s travelling direction, mathematically given through the dot product between the two normalized vectors (thresholding it at an angle of 10). Within this pool of filtered edges, we choose a source node which fulfills the property of being closer to $FT$; found by calculating the L1-norm between $s_{i}^{0}$ and the nearest nodes.
A depth-first search (DFS) is performed on each of the nearest nodes to find the goal lanes. For each goal lane, the algorithm checks if it is in an intersection. If any one of the lanes $l$ in $L_{goal}$ is present in an intersection, the trajectory is returned with a scenario tag of 1; otherwise 0. This process allows for the categorization of scenes into two bins based on the presence of goal lanes in an intersection, as can be seen from algorithm 1.
Algorithm 1 Scene-Specific Set Generation
1:Vector-Map $M$, State Information of Focal Track $FT$
2:for each Scene do
3:     Construct a directed graph $G$ from $M$
4:     $S_{past}$, $\theta$leftarrow Query for the focal track, where $\theta$ is the heading angle
5:     $R\leftarrow$ Calculate Rotation Matrix($\theta$)
6:     $N_{nearest}\leftarrow$ Filter for edges with parallel direction vectors to $FT$
7:     $n_{source}\leftarrow$ Find the source node using L1-norm between $S_{t}[-1]$ and $N_{nearest}$
8:     $L_{goal}\leftarrow$ Perform DFS on all the nodes in $N_{nearest}$
9:     for $l$ in $L_{goal}$ do
10:         if $l$ is in an Intersection then
11:              return Trajectory, $T_{scene}=1$
12:         else
13:              return Trajectory, $T_{scene}=0$
14:         end if
15:     end for
16:     return None
17:end for
Running algorithm 1 across the training dataset yields two different candidate sets that reflect the local representation of the scene, as shown in fig. 3 and fig. 3. The trajectories corresponding to the scenes are normalized to $FT$’s last observation state using the rotation matrix $R$ from line 4 of algorithm 1. This approach maximizes the probability of finding trajectories closer to the ground truth without compromising on the shortcomings of earlier approaches which utilize a single set. To optimize the candidate sets further, we need to determine two key parameters: the optimal set size and the sampling strategy. The sampling approach plays a crucial role in reducing the candidate sets to an optimal size suitable for the intended application.
V Metrics
The distance between trajectories within a set is quantified using Euclidean metrics, such as the minimum Average Displacement Error (minADE) and the minimum Final Displacement Error (minFDE). A diverse set of trajectories serves as a representative sample of the underlying population, capturing its intrinsic characteristics. By assessing the properties of each trajectory within the set, we can quantify its individual contribution to the overall diversity. To measure the plausibility of the set with respect to the drivable regions of a scene, we employ the Driving Area Compliance (DAC) score, which measures the compatibility of a set for a given scene. Therefore,
$DAC=\frac{\kappa-n}{\kappa}$
where, $\kappa$ represents the given set size and $n$ denotes the number of non-compliant trajectories which goes off-road. A higher DAC score provides an indication of the set’s density in relation to the scene. We use the same methodology as described in [5] to get the DAC value per scene.
By integrating the DAC and Euclidean metrics, we enhance our assessment of the set’s quality, which hinges on its closeness to the ground truth trajectories. A higher DAC score signifies that more trajectories adhere to drivable areas, whereas lower minADE and minFDE values indicate closer alignment with the ground truth. Employing both diversity and scene compliance metrics ensures that the optimized set of trajectories maximizes the representativeness of the set, providing a comprehensive portrayal of the true population dynamics w.r.t the overall map geometry.
VI Experiments
We use the motion forecasting dataset from Argoverse 2 (AV2)[21], which contains 199,489 training samples or scenes and 20,000 validation samples. The construction and analysis of the validation samples are refrained from, as that would lead to look-ahead bias. Therefore, all the analysis is performed on the training distribution unless explicitly stated otherwise.
The dataset contains HD maps as a lane graph, from which we construct a directed Graph similar to VectorNet polyline [22] representation. Every scene has a target actor called as a focal track ($FT$), whose observation window satisfies the requirement of possessing full state information throughout the complete sampling horizon. For AV2, the sampling horizon is 6s for the future, with a 5s observation history; sampled at 10Hz. A Rotation matrix $R$ is calculated using the heading angle and the last XY coordinate at a time step $t_{\text{obs}}$. The trajectory $T_{i}$ is transformed into a local coordinate system (actor-centric) by applying $R$, which is better suited to represent trajectories.
TABLE I: Comparison of coverage-based approach vs. random sampling from the training distribution.
Method Set Size LB. minADE ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ )
1k 0.63 0.525
1.5k 0.58 0.520
2k 0.55 0.499
Metric Driven 2.5k 0.52 0.494
3k 0.51 0.490
4k 0.48 0.481
5k 0.45 0.486
1k (854) 0.771 (0.782) 0.611 (0.673)
1.5k (1246) 0.710 (0.731) 0.627 (0.662)
2k (1642) 0.661 (0.667) 0.607 (0.664)
Ran-Samp(0.2m) 2.5k (2067) 0.617 (0.621) 0.615 (0.678)
3k (2450) 0.598 (0.603) 0.608 (0.682)
4k (3237) 0.559 (0.564) 0.606 (0.671)
5k (4023) 0.532 (0.538) 0.634 (0.675)
A trajectory set for a training distribution $T_{full}$ is constructed by parsing through each scene and querying for the focal track’s future position. Since each of the trajectory is normalized with their respective $R$, we get a set which resembles fig. 1.
VI-A Evaluation of a single set
Validating the effectiveness of the sets is crucial to comprehend the influence of two primary parameters which govern diversity and plausibility. We employ the metric-driven method [12] and a random sampling to demonstrate the efficacy of our proposed scene-specific set generation concept. Similar to the Covernet [3]; metric-driven method also employs a minADE driven bagging algorithm to reduce each of the candidate sets, we have acquired from the algorithm 1. Random sampling (Ran-Samp) is used as a baseline, and we report the mean values as it exhibits a variance of ±$\pm\text{2\%}$%.
We compute a theoretical Lower Bound (LB) which measures the quality of the constructed set validated against the trajectories from the training distribution. This lower bound is derived from the minADE between the ground truth trajectory and the closest trajectory within the set. The proposed approach enables us to assess the quality and diversity of the generated trajectory sets. By comparing the performance of our method against the baseline and analyzing the lower bounds, we gain insights into how effective our set construction is in capturing diverse and plausible trajectories. table I illustrates the effect of metric-driven sets when compared with randomly sampled sets generated from the complete training distribution.
As the set size increases, the LB of the metrics linearly decreases, this is attributed to the presence of more trajectories which lie closer to the GT trajectory, thereby providing good coverage. Sets generated through the metric-driven reduction method perform well as expected, since the cost function enforces diversity between the trajectories in the set when compared with Ran-Samp.
On the other hand, from fig. 5 emerges a wholly different picture. The normalized rate of surviving trajectories, that which are compliant with the scene, is substantially higher for the Ran-Samp and amounts to $10\%$ more trajectories irrespective of the set size.
The Ran-Samp method, however, is not without its drawbacks. It includes trajectories that overlap and are deemed redundant, which could be eliminated without adversely affecting the metrics. To address this, we introduce a thresholding rate of $0.2$m, enabling the removal of redundant trajectories that lie close to each other within the Ran-Samp set. This refinement process results in a more lean set, reducing the number of trajectories by approximately $15\%$ compared to the original set size, as can be seen from table I. Moreover, this process improves the DAC rate while preserving an equivalent minADE when juxtaposed with simple Ran-Samp set.
Figure 5: The randomly sampled sets can retain more number of trajectories when compared with a metric-driven set, making the former a reliable way to represent more plausible states
VI-B Evaluation of scene-specific sets
Utilizing algorithm 1, we generated two distinct candidate sets: one pertaining to intersections and the other to non-intersections. Subsequently, for each candidate set, adhering to their corresponding sampling criteria, we performed extensive analysis, as can be seen from the table II.
The LB.minADE column of table II, when evaluated against the results from the table I shows similar correlation between the internal representation of trajectories (sets’ diversity) and the set size. As we incrementally increase the set size from 1k to 5k, the LB.minADE shows a consistent decrease for both methods (metric-driven and Ran-Samp) and across both scenario types (Non-intersection and intersection set). This suggests that the generation of scene-specific sets increases the probability of finding more trajectories that lie closer to the ground truth.
TABLE II: Quantitative comparison of Scene-specific sets with different methods on Argoverse training distribution
Non-Intersection Set Intersection Set
Method Set Size LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ set ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ ) LB. minADE ()($\downarrow$)( ↓ ) LB. minADE w/ set ()($\downarrow$)( ↓ ) DAC()($\uparrow$)( ↑ )
1k 0.59 0.67 0.694 0.72 0.73 0.518
1.5k 0.53 0.62 0.673 0.66 0.67 0.500
2k 0.50 0.59 0.677 0.62 0.63 0.490
Metric Driven 2.5k 0.47 0.56 0.664 0.59 0.60 0.484
3k 0.44 0.54 0.704 0.57 0.58 0.477
4k 0.41 0.52 0.692 0.54 0.55 0.468
5k 0.38 0.49 0.687 0.51 0.52 0.469
1k 0.72 0.83 0.827 0.874 0.89 0.691
1.5k 0.64 0.74 0.831 0.789 0.81 0.664
2k 0.60 0.68 0.823 0.729 0.74 0.663
Ran-Samp 2.5k 0.56 0.65 0.842 0.690 0.70 0.678
3k 0.54 0.63 0.833 0.659 0.68 0.692
4k 0.49 0.59 0.835 0.606 0.62 0.681
5k 0.45 0.55 0.863 0.579 0.58 0.679
1k 0.712 0.824 0.866 0.689
1.5k 0.661 0.827 0.779 0.658
2k 0.591 0.826 0.717 0.656
RIDS 2.5k 0.557 0.836 0.676 0.650
3k 0.532 0.829 0.651 0.682
4k 0.488 0.839 0.598 0.688
5k 0.455 0.850 0.572 0.681
The intuition behind reporting LB.minADE w/ set, is to establish a baseline to compare how a method (metric-driven, ran-samp, RIDS); pertaining to a specific set constructed from a full training distribution, would fare against different scenes (non-intersection/intersection). This approach allows for a fair comparison between different set construction methods and helps identify the most suitable method for specific types of scenes.
The DAC metric (normalized surviving trajectories per scene) exhibits relative stability across different set sizes for each method and scenario combination. The Ran-Samp method consistently shows higher DAC values compared to metric-driven, suggesting that it is more effective in generating a set where a larger proportion of the trajectories remain plausible. When comparing the two scenario types, it is evident that both methods yield poorer performance in intersection scenarios compared to non-intersections, as measured between corresponding LB.minADE w/ set and LB.minADE metrics. This outcome is anticipated due to the increased complexity and diversity of potential trajectories at intersections. However, the performance degradation from non-intersection to intersection scenarios is less evident for Ran-Samp than for metric-driven. As set sizes increase, the displacement error narrows, and the number of surviving trajectories experiences a significant increase, a trend that is also observable in the intersection sets.
We also created a new subsampling algorithm, RIDS or Recursive In Distribution Subsampling, which is a method where we remove duplicate trajectories from a set constructed using Ran-samp. If two trajectories share proximity to each other and lie within a threshold of 0.2m to each other, we consider this trajectory redundant and arbitrarily remove one of them. This leaves us with a set which is underweighted; to counter this, we perform subsampling from the training distribution and recursively perform this step of removing and adding trajectories until the desired $T^{\prime}$ is reached.
VII Results and Discussions
The ultimate goal of this work is to find the answer to the following question: what is the best conceptual framework that can maximize the probability of finding trajectories closer to the ground truth? This led us to investigate and propose a scene-specific set generation method, which, when viewed through the lenses of displacement errors and DAC, gives better performance compared to an uninformed set. The next step is to find the sampling method; from fig. 5, random sampling provides preliminary evidence, as it retains more trajectories than the metric-driven when one wants to achieve a dense set representation.
Trajectory sets are used for the downstream task of motion prediction of actors in the environment, where the probability of finding more trajectories outweighs the minimal performance gains from the displacement metrics. In this case, a set that maximizes plausibility is desired if the nominal gains from the Euclidean metrics are minimal, since we can guarantee that there exist more trajectories in the proximity which can represent the ground truth.
Although the metric-driven method was able to provide guarantees for probabilistic completeness for maximizing coverage, it failed to capture the true distribution, providing insufficient exploration. On the other hand, sampling from a uniform distribution; RIDS was able to retain more trajectories and outperform the metric-driven method. The property of an efficient estimator is one that has the smallest possible variance, indicating that it is closest to the true value of the parameter being estimated. This reasoning that RIDS being an efficient estimator and functions effectively because in certain optimization situations, sampling from the underlying distribution can be more efficient than performing sampling based on some metrics.
Our experiments have shown that the scene-specific set generation method allows us to improve diversity, and more specifically, RIDS as a sampling method works relatively well when considering the environment, resulting in an improved DAC score. The survival rate for trajectories for sets constructed using Ran-Samp and RIDS outperforms metric-driven sampling methods, given that the minimal gain in displacement errors from metric-driven overshadows the presence of more plausible trajectories. A visual distinction is notable for the non-intersection set in fig. 3, which confirms with our prior, as most of the trajectories represent no significant yaw differences (going straight). The DAC score for the non-intersection with Ran-Samp depicts that cluster-based trajectory sampling works better than expected, since it can retain 86% of the trajectories for $T^{5k}_{set}$. This high score indicates that the set can handle the complex geometry of the real-world and is perfectly suited for the intended application. As we said at the start, this showcases that Scene-Specific Set generation can produce sets which can be tuned to represent the ground truth closer to real-world scenes than using a single set to represent diverse situations.
Looking at Ran-Samp set created with a threshold rate of 0.2 from table I; almost one could say that this imbalance between the set size is the reason for the improved DAC value. To counter that argument, we performed ablation studies to confirm that these improved DAC values remain improved when performing the RIDS sampling process for different threshold values.
VIII Conclusion
In this work, we presented our approach to create scene specific set generation which retains more trajectories akin to the environment which aims to maximize for the plausibility of trajectories for a diverse amount of environment. Along with that, we also showcased that RIDS as a sampling process is effective in condensing the full representation space to a more manageable set while providing coverage and plausibility to the corresponding set.
Acknowledgment
The research leading to these results was funded by the German Federal Ministry for Economic Affairs and Climate Action and was partially conducted in the project “KI Wissen”. We would also like to thank Max Zipfl and Ahmed Abouelazm for their valuable reviews. Responsibility for the information and views set out in this publication lies entirely with the authors.